OpenStudy (anonymous):

There is a formula used in chemistry and physics that states that PV = nrt. Solve this formula for n.

OpenStudy (anonymous):

Welcome to open study!!

OpenStudy (bibby):

PV=n*r*t or \(PV=n\times r \times t \) if multiplication and division are inverse operations how would you isolate n?